OpenStudy (anonymous):

Solve for x. -3x - 8 + 6x = 7

OpenStudy (anonymous):

x=-7/5 x=-1/3 x=5 x=45

OpenStudy (anonymous):

@phi

OpenStudy (kirbykirby):

-3x - 8 + 6x = 7 -3x - 8 = 7 -3x = 7 + 8 -3x = 15 x = 15/(-3) x = -5

OpenStudy (anonymous):

Wait so c